OpenStudy (anonymous):

which of the following statements are true about the graph of y=ln(4+x^2) a)it is symmetric on the y axis b)it has a local minimum at x=0 c) it has inflection points at = +/- 2

ganeshie8 (ganeshie8):

hey show your work

OpenStudy (anonymous):

what do u mean? I don't know how to solve it..

ganeshie8 (ganeshie8):

A graph is symmetric about y axis if `f(x) = f(-x)`

ganeshie8 (ganeshie8):

you may use that for part a

OpenStudy (anonymous):

ok but how would i find that out with that function?

ganeshie8 (ganeshie8):

\(y = f(x) = \ln(4+x^2)\) replace x by -x for f(-x) : \(f(-x) = \ln(4+(-x)^2) = \ln(4+x^2) = f(x)\) that means...

OpenStudy (anonymous):

that it isn't symmetric on the y axis

ganeshie8 (ganeshie8):

think again

OpenStudy (anonymous):

ohh that it is symmetrical

ganeshie8 (ganeshie8):

Yep! thats it for part a see if you can work remaining parts using first/second derivatives

OpenStudy (anonymous):

thanks but I'm not sure how to get the first and second derivative since there is an ln..

ganeshie8 (ganeshie8):

use the derivative of ln and chain rule

ganeshie8 (ganeshie8):

\[\dfrac{d}{dx}(\ln x) = \frac{1}{x}\]

OpenStudy (usukidoll):

problem is on the second derivative, it's gonna get nastay

OpenStudy (usukidoll):

ok probably not.. I accidentally left the ln :P/ but second derivative is quotient rule

OpenStudy (anonymous):

wait why would i use the chain rule

OpenStudy (usukidoll):

hmm there's no exponents I don't see the point.. umm do you know how to take derivatives for ln?

OpenStudy (usukidoll):

basically it's 1/ whatever it is inside the ln x (derivative of ln) so for example I have ln(2x)=y 1/2x (2) = y'

OpenStudy (usukidoll):

which simplifies to 1/x = y' second time you take the derivative.. quotient rule

OpenStudy (usukidoll):

you know I have no idea why I can do math maturely on here... but when I'm in the class with this one particular guy prof.. quick I can't get my mind to focus.

OpenStudy (anonymous):

wait the derivative of ln is 1/x, and then the derivative of (4+x^2) is 2x so we're left with 1/x(2x) right ?

OpenStudy (usukidoll):

ummmmmmm that's just ln x default we're given ln 4+x^2 yes the derivative is 2x but for derivative of ln we need 1 / (original function ) x (derivative inside the original function) 1/(4+x^2) x (2x)

OpenStudy (usukidoll):



OpenStudy (usukidoll):

then it's quotient rule on that guy f/g gf'-fg' / g^2

OpenStudy (anonymous):

oh ok i see. so i must take the second derivative?

OpenStudy (usukidoll):

if we want to know concavity .. ? hmm it's not asking that just what interval are we in?

OpenStudy (anonymous):

no it's just asking to know what is true about the graph. i have the answer choices on the question.

OpenStudy (anonymous):

it could be all of them or 2 of them or only one..